# DynaSpec: Context-aware Dynamic Speculative Sampling for Large-Vocabulary Language Models

### v1

@@ Abstract
Speculative decoding (a.k.a. speculative sampling) has become a standard way to accelerate LLM inference: a small drafter proposes multiple tokens and a large target model verifies them once per speculation length. Recently, scaling of the LLM vocabulary has pushed the number of tokens to grow substantially. While verification over the full vocabulary leaves the target model largely unaffected, the $\mathcal{O}(|V|d)$ parameters in the drafter’s output head become a latency bottleneck, slowing the entire pipeline. Contemporary methods (e.g., FR-Spec, VocabTrim) restrict the drafter’s vocabulary to a fixed subset of the target model’s vocabulary, ranked in descending order of token frequency. Although this reduces draft-time compute, it is brittle, since : (i) frequency lists are corpus-dependent and require retuning to generalize, and (ii) static shortlists suppress rare or domain-specific tokens, lowering the expected number of tokens per verification step. We propose DynaSpec, a context-dependent dynamic shortlisting mechanism that is robust, speeds up drafting, and generalizes across diverse tasks. Concretely, we introduce lightweight, coarse-grained meta-classifiers that route contexts to a small number of token clusters; the union of the top-$k$ selected clusters forms the drafter’s shortlist, while verification retains the full vocabulary and exactness. The meta-classifier finishes its computation earlier than the drafter’s hidden state generation by exploiting parallel execution of draft encoding and meta shortlisting on separate streams. On standard speculative-decoding benchmarks, we observe consistent gains in mean accepted length over fixed-shortlist baselines, while context-dependent selection enables smaller shortlists without degrading acceptance.
@@ 1 Introduction
Over the past few years, large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Liu et al., 2024; Jiang et al., 2024) have improved dramatically, reshaping products and workflows across sectors. Beyond algorithmic advances, much of this progress has come from scaling data and especially model parameters (Kaplan et al., 2020; Hoffmann et al., 2022). Yet these gains come with a cost: inference becomes compute-heavy and latency-sensitive, which limits real-time applications (Kwon et al., 2023; Yuan et al., 2024). As agentic systems (Yao et al., 2023; Wang et al., 2024) and reasoning-heavy workflows (e.g., chain-of-thought) (Wei et al., 2022; Zhang et al., 2024) proliferate, the demand for faster inference keeps rising. Speculative decoding (Chen et al., 2023; Leviathan et al., 2023; Li et al., 2024a, b) is a practical way to accelerate generation. A smaller draft model proposes multiple tokens so the target model need not run at every step, while the rejection sampling procedure guarantees that the final output distribution matches that of the target model. In the worst case, quality and latency revert to standard decoding; in typical cases, throughput improves substantially.
@@
Recently scaling laws for vocabulary (Tao et al., 2024; Huang et al., ; Yu et al., 2025) suggests that larger models benefit from larger vocabularies, and deployed LLMs have indeed expanded their vocabularies over time. For example, tokenizer vocabularies have scaled from 32k entries in Llama-2 (Touvron et al., 2023) to 128k in Llama-3 (Grattafiori et al., 2024), 129k in DeepSeek-V3 (Liu et al., 2024), 152k in Qwen-2.5 (Yang et al., 2025), and 262k in Gemma-3 (Team et al., 2025). While the large vocabulary size contributes only marginally to the inference cost of the target model’s output layers given its overall model size, it imposes a substantial computational burden on the draft model during speculative decoding, as the draft model is much smaller and the output layer constitutes a larger fraction of its total computation. thereby diminishing the overall pipeline efficiency. To mitigate this, FR-Spec (Zhao et al., 2025) and Vocabtrim (Goel et al., 2025) replace the full vocabulary with a fixed shortlist—typically the top-p% of the most frequent tokens measured on a reference corpus. This speeds up drafting, but it is suboptimal for two reasons: (1) frequency lists are corpus-dependent and often fail to generalize across benchmarks, requiring substantial retuning; and (2) a static subset of the token vocabulary can degrade performance on tasks that rely on rare or domain-specific tokens, where maintaining diversity in the candidate set is critical for high-quality outputs.
@@
In this work, we move beyond fixed shortlists and introduce a dynamic vocabulary head. We first define meta-labels as coarse token clusters: the vocabulary is partitioned into $M\!\ll\!|V|$ clusters $\{C_{m}\}_{m=1}^{M}$ by clustering column-normalized LM-head weights. Conditioned on embedding and features, a lightweight meta-classifier scores clusters and we evaluate the drafter’s LM head only on the union of the selected clusters $\mathcal{V}_{S}=\bigcup_{m\in\mathcal{K}}\!C_{m}$, while verification still runs over the full vocabulary, preserving exactness. This coarse-to-fine routing, inspired by Jiang et al. (2021) and  Kharbanda et al. (2022), replaces an $\mathcal{O}(|V|d)$ head with $\mathcal{O}\!\big((M+|\mathcal{V}_{S}|)d\big)$. Making the support context-dependent leads to a consistent increase (across datasets) in the mean accepted length relative to the static subset proposed in Zhao et al. (2025). However, a dynamic (indexed) head introduces gathered matmul overhead and can be slower than a fixed, pre-trimmed dense head. We therefore couple the dynamic head with a position-aware cluster budget: early tokens receive larger shortlists, and the budget decays with position. As argued in recent works, Li et al. (2025, 2024b), prioritizing earlier tokens improves expected acceptance, while shrinking the shortlist later reduces latency, yielding net end-to-end gains without altering verification. Finally, we show the same position-aware schedule can retrofit frequency-ranked shortlists (FR-Spec) (Zhao et al., 2025); see Appendix A.1 for details.
@@ 2.1 LLMs with Large Vocabulary
As LLMs have grown in scale, so too have their vocabularies. Early models such as GPT-3 (Brown et al., 2020) and LLaMA-2 (Touvron et al., 2023) used vocabularies of $32\mathrm{k}$ tokens, while more recent releases exceed $100\mathrm{k}$ (e.g., $128\mathrm{k}$ in LLaMA-3 (Grattafiori et al., 2024), $152\mathrm{k}$ in Qwen-2.5 (Yang et al., 2025), and $262\mathrm{k}$ in Gemma-3 (Team et al., 2025)). Large vocabularies are especially common in multilingual and high-coverage LLMs: mT5 (Xue et al., 2020) adopts a $250\mathrm{k}$ SentencePiece (Kudo and Richardson, 2018) vocabulary to improve cross-lingual coverage, and PaLM uses a $256\mathrm{k}$ unigram vocabulary to shorten sequences and better encode code and symbols (Chowdhery et al., 2023). Recently, empirical scaling laws suggest that larger vocabularies improve expressivity and perplexity (Tao et al., 2024; Yu et al., 2025), but they also magnify inference cost: the output projection grows linearly with vocabulary size, which can dominate latency in small drafters.
@@ 2.2 Speculative Decoding
In line with Speculative Decoding, early work accelerates greedy generation (Stern et al., 2018; Sun et al., 2021), and subsequent work extends the idea to non-greedy sampling with distributional guarantees (Leviathan et al., 2023; Chen et al., 2023; Xia et al., 2022). These methods differ primarily in how they form proposals: some reuse information already present in the prompt via retrieval-guided drafting—effective when outputs strongly overlap inputs (Saxena, 2023; Yang et al., 2023; He et al., 2024)—whereas others learn parametric drafters for general-purpose speedups, ranging from parallel MLP heads (and serial variants) to lightweight transformer blocks that condition on richer context (Cai et al., 2024; Ankner et al., 2024). The EAGLE series further refines this paradigm: EAGLE employs lightweight transformer drafters, while EAGLE-2 introduces dynamic draft trees to boost efficiency (Li et al., 2024a, b). Orthogonal advances in training and systems—such as adaptive exploration, cache reuse, and efficient serving stacks—increase accepted tokens per verification and reduce overhead (Kwon et al., 2023; Zheng et al., 2024; Dao, 2023). Complementing these directions, MagicDec analyzes the latency–throughput trade-off and shows that speculative decoding remains effective at large batch sizes and with long contexts, whereas TriForce targets long-context regimes by compressing the drafter’s KV cache and coordinating reuse across draft steps (Sadhukhan et al., 2024; Sun et al., 2024). Most recently, Timor et al. (2025) proposes lossless speculative decoding for heterogeneous vocabularies, relaxing the same-tokenizer requirement between drafter and target; meanwhile, Li et al. (2025) combines Medusa-style heads (Cai et al., 2024) with the EAGLE (Li et al., 2024b) framework. We target a different bottleneck: as vocabularies expand, the drafter’s output projection increasingly dominates latency.
@@ 2.3 Accelerating Large Vocabularies
The computational bottleneck of large output spaces or vocabularies has been recognized in both language modeling and neighboring fields. Within LLMs, FR-Spec (Zhao et al., 2025) and VocabTrim (Goel et al., 2025) mitigate drafter overhead by pruning the output layer to a fixed shortlist of high-frequency tokens. While effective in reducing latency, these static subsets may suppress rare or domain-specific tokens, lowering acceptance in speculative decoding. Other work on large vocabulary LLM training, such as CCE (Wijmans et al., 2024), reduces peak memory by fusing cross-entropy directly into the kernel, avoiding explicit logit materialization. Our approach is motivated by research in extreme classification, where the goal is to accelerate learning and inference over output spaces with millions of labels. Methods such as LightXML (Jiang et al., 2021), CascadeXML (Kharbanda et al., 2022), and related meta classifier based strategies (Zhang et al., 2021; Kharbanda et al., 2023) adopt coarse-to-fine or cluster-based mechanisms that limit computation to a small subset of relevant labels per input. These ideas directly inspire DynaSpec: rather than evaluating the drafter across the entire vocabulary or a static shortlist, we employ a lightweight meta-classifier to select a small number of context-relevant clusters, retaining the coverage of rare tokens while reducing compute.
@@ 3 Theoretical Analysis
Let $T_{T}$ denote the per-token latency of the target model under standard autoregressive decoding. In speculative decoding, let $T_{D}$ be the draft model’s per-token generation time, and let $T_{V}(\gamma)$ be the time for the target to verify $\gamma$ proposed tokens. Given $\alpha\in[0,1]$ as the expected acceptance rate and $\gamma\in\mathbb{N}$ the speculation length, following standard derivations (Leviathan et al., 2023), the expected number of tokens produced per verification step is
@@ 3.2 Impact of Static Sub-Vocabulary Constraints on Acceptance
 [^1] For details, follow section 3 in (Leviathan et al., 2023)
@@
$\beta(c;q_{c})\;=\;\sum_{x\in V}\min(p_{c}(x),\,q_{c}(x))\;=\;\mathbb{E}_{x\sim q_{c}}\!\left[\min\!\Big(1,\tfrac{p_{c}(x)}{q_{c}(x)}\Big)\right].$ (3)
@@
$A^{\!*}(c;S)\;:=\;\max_{q_{c}:\ \mathrm{supp}(q_{c})\subseteq S}\ \sum_{x\in V}\min\{p_{c}(x),q_{c}(x)\}\;=\;\sum_{x\in S}p_{c}(x)\,$
@@
$\mathbb{E}_{c}\!\big[A^{\!*}(c;S_{c}^{\star})\big]\;=\;\mathbb{E}_{c}\!\big[H_{k}(p_{c})\big]\;\geq\;\mathbb{E}_{c}\!\big[A^{\!*}(c;S_{\mathrm{stat}})\big]\;=\;\mathbb{E}_{c}\!\Big[\sum_{x\in S_{\mathrm{stat}}}p_{c}(x)\Big],$
@@ 4.1 Preliminary
Full Vocabulary Drafter. Speculative decoding methods (more specifically EAGLE (Li et al., 2024a, b) series of models) train a lightweight draft model $\mathcal{D}$ (one transformer block, same latent width $d$) to approximate $\mathcal{T}$ at a lower cost. $\mathcal{D}$ reuses and freezes the embedding $\mathbf{E}$ and LM head $W_{LM}$ from $\mathcal{T}$; only its single transformer block is trained:
@@
Static Sub-vocabulary Drafter. FR-Spec (Zhao et al., 2025) uses a frequency-ranked drafting with vocabulary $\mathcal{V}_{\text{high}}\!\subset\!\mathcal{V}$ the subset of high-frequency tokens (from corpus statistics). Concretely, select the corresponding LM-head columns.
@@ 4.2 Meta classifiers: Context-aware token Selection
Instead of a fixed top-$k$ policy, we use a position-aware budget $k_{\mathrm{c}}(t)\in\{k_{\min},\ldots,k_{\max}\}$ that allocates more clusters to early tokens and fewer to later ones. Concretely, with $t=0$ denoting the first drafted token, we keep the first two steps at the full budget and then decay as
$k_{\mathrm{c}}(t)=\begin{cases}k_{\max},&t\in\{0,1\},\\[4.0pt] \left\lfloor\dfrac{k_{\max}}{(t+1)\cdot 2}\right\rfloor,&t\geq 2.\end{cases}$
Given router scores $s$, we select $\mathcal{K}(c,t)=\operatorname{TopK}_{k_{\mathrm{c}}(t)}(s)$ and construct the shortlist $\mathcal{V}_{S}(c,t)=\bigcup_{m\in\mathcal{K}(c,t)}C_{m}$.
@@
$\mathcal{L}_{\mathrm{router}}\;=\;\mathbb{E}\!\left[\sum_{m=1}^{M}\Big(-\,y_{m}\log\hat{y}_{m}-(1-y_{m})\log\big(1-\hat{y}_{m}\big)\Big)\right].$

### v2

@@ Abstract
[…]
@@ 1 Introduction
Over the past few years, large language models (LLMs) (brown2020language; touvron2023llama; liu2024deepseek; jiang2024mixtral) have improved dramatically, reshaping products and workflows across sectors. Beyond algorithmic advances, much of this progress has come from scaling data and especially model parameters (kaplan2020scaling; hoffmann2022training). Yet these gains come with a cost: inference becomes compute-heavy and latency-sensitive, which limits real-time applications (kwon2023efficient; yuan2024llm). As agentic systems (yao2023react; wang2024survey) and reasoning-heavy workflows (e.g., chain-of-thought) (wei2022chain; zhang2024chain) proliferate, the demand for faster inference keeps rising. Speculative decoding (chen2023accelerating; leviathan2023fast; li2024eagle; li2024eagle2) is a practical way to accelerate generation. A smaller draft model proposes multiple tokens so the target model need not run at every step, while the rejection sampling procedure guarantees that the final output distribution matches that of the target model. In the worst case, quality and latency revert to standard decoding; in typical cases, throughput improves substantially.
@@
Recently scaling laws for vocabulary (tao2024scaling; huangover; yu2025scaling) suggests that larger models benefit from larger vocabularies, and deployed LLMs have indeed expanded their vocabularies over time. For example, tokenizer vocabularies have scaled from 32k entries in Llama-2 (touvron2023llama) to 128k in Llama-3 (grattafiori2024llama), 129k in DeepSeek-V3 (liu2024deepseek), 152k in Qwen-2.5 (yang2025qwen3), and 262k in Gemma-3 (team2025gemma). While the large vocabulary size contributes only marginally to the inference cost of the target model’s output layers given its overall model size, it imposes a substantial computational burden on the draft model during speculative decoding, as the draft model is much smaller and the output layer constitutes a larger fraction of its total computation. thereby diminishing the overall pipeline efficiency. To mitigate this, FR-Spec (zhao-etal-2025-fr) and Vocabtrim (goel2025vocabtrim) replace the full vocabulary with a fixed shortlist—typically the top-p% of the most frequent tokens measured on a reference corpus. This speeds up drafting, but it is suboptimal for two reasons: (1) frequency lists are corpus-dependent and often fail to generalize across benchmarks, requiring substantial retuning; and (2) a static subset of the token vocabulary can degrade performance on tasks that rely on rare or domain-specific tokens, where maintaining diversity in the candidate set is critical for high-quality outputs.
@@
In this work, we move beyond fixed shortlists and introduce a dynamic vocabulary head. We first define meta-labels as coarse token clusters: the vocabulary is partitioned into $M\!\ll\!|V|$ clusters $\{C_{m}\}_{m=1}^{M}$ by clustering column-normalized LM-head weights. Conditioned on embedding and features, a lightweight meta-classifier scores clusters and we evaluate the drafter’s LM head only on the union of the selected clusters $\mathcal{V}_{S}=\bigcup_{m\in\mathcal{K}}\!C_{m}$, while verification still runs over the full vocabulary, preserving exactness. This coarse-to-fine routing, inspired by jiang2021lightxml and  kharbanda2022cascadexml, replaces an $\mathcal{O}(|V|d)$ head with $\mathcal{O}\!\big((M+|\mathcal{V}_{S}|)d\big)$. Making the support context-dependent leads to a consistent increase (across datasets) in the mean accepted length relative to the static subset proposed in zhao-etal-2025-fr. However, a dynamic (indexed) head introduces gathered matmul overhead and can be slower than a fixed, pre-trimmed dense head. We therefore couple the dynamic head with a position-aware cluster budget: early tokens receive larger shortlists, and the budget decays with position. As argued in recent works, li2025gumiho; li2024eagle2, prioritizing earlier tokens improves expected acceptance, while shrinking the shortlist later reduces latency, yielding net end-to-end gains without altering verification. Finally, we show the same position-aware schedule can retrofit frequency-ranked shortlists (FR-Spec) (zhao-etal-2025-fr); see Appendix A.1 for details.
@@ 2.1 LLMs with Large Vocabulary
As LLMs have grown in scale, so too have their vocabularies. Early models such as GPT-3 (brown2020language) and LLaMA-2 (touvron2023llama) used vocabularies of $32\mathrm{k}$ tokens, while more recent releases exceed $100\mathrm{k}$ (e.g., $128\mathrm{k}$ in LLaMA-3 (grattafiori2024llama), $152\mathrm{k}$ in Qwen-2.5 (yang2025qwen3), and $262\mathrm{k}$ in Gemma-3 (team2025gemma)). Large vocabularies are especially common in multilingual and high-coverage LLMs: mT5 (xue2020mt5) adopts a $250\mathrm{k}$ SentencePiece (kudo2018sentencepiece) vocabulary to improve cross-lingual coverage, and PaLM uses a $256\mathrm{k}$ unigram vocabulary to shorten sequences and better encode code and symbols (chowdhery2023palm). Recently, empirical scaling laws suggest that larger vocabularies improve expressivity and perplexity (tao2024scaling; yu2025scaling), but they also magnify inference cost: the output projection grows linearly with vocabulary size, which can dominate latency in small drafters.
@@ 2.2 Speculative Decoding
In line with Speculative Decoding, early work accelerates greedy generation (stern2018blockwise; sun2021instantaneous), and subsequent work extends the idea to non-greedy sampling with distributional guarantees (leviathan2023fast; chen2023accelerating; xia2022speculative). These methods differ primarily in how they form proposals: some reuse information already present in the prompt via retrieval-guided drafting—effective when outputs strongly overlap inputs (saxena2023prompt; yang2023inference; he2024rest)—whereas others learn parametric drafters for general-purpose speedups, ranging from parallel MLP heads (and serial variants) to lightweight transformer blocks that condition on richer context (cai2024medusa; ankner2024hydra). The EAGLE series further refines this paradigm: EAGLE employs lightweight transformer drafters, while EAGLE-2 introduces dynamic draft trees to boost efficiency (li2024eagle; li2024eagle2). Orthogonal advances in training and systems—such as adaptive exploration, cache reuse, and efficient serving stacks—increase accepted tokens per verification and reduce overhead (kwon2023efficient; zheng2024sglang; dao2023flashattention). Complementing these directions, MagicDec analyzes the latency–throughput trade-off and shows that speculative decoding remains effective at large batch sizes and with long contexts, whereas TriForce targets long-context regimes by compressing the drafter’s KV cache and coordinating reuse across draft steps (sadhukhan2024magicdec; sun2024triforce). Most recently, timor2025accelerating proposes lossless speculative decoding for heterogeneous vocabularies, relaxing the same-tokenizer requirement between drafter and target; meanwhile, li2025gumiho combines Medusa-style heads (cai2024medusa) with the EAGLE (li2024eagle2) framework. We target a different bottleneck: as vocabularies expand, the drafter’s output projection increasingly dominates latency.
@@ 2.3 Accelerating Large Vocabularies
The computational bottleneck of large output spaces or vocabularies has been recognized in both language modeling and neighboring fields. Within LLMs, FR-Spec (zhao-etal-2025-fr) and VocabTrim (goel2025vocabtrim) mitigate drafter overhead by pruning the output layer to a fixed shortlist of high-frequency tokens. While effective in reducing latency, these static subsets may suppress rare or domain-specific tokens, lowering acceptance in speculative decoding. Other work on large vocabulary LLM training, such as CCE (wijmans2024cut), reduces peak memory by fusing cross-entropy directly into the kernel, avoiding explicit logit materialization. Our approach is motivated by research in extreme classification, where the goal is to accelerate learning and inference over output spaces with millions of labels. Methods such as LightXML (jiang2021lightxml), CascadeXML (kharbanda2022cascadexml), and related meta classifier based strategies (zhang2021fast; kharbanda2023inceptionxml) adopt coarse-to-fine or cluster-based mechanisms that limit computation to a small subset of relevant labels per input. These ideas directly inspire DynaSpec: rather than evaluating the drafter across the entire vocabulary or a static shortlist, we employ a lightweight meta-classifier to select a small number of context-relevant clusters, retaining the coverage of rare tokens while reducing compute.
@@ 3 Theoretical Analysis
Let $T_{T}$ denote the per-token latency of the target model under standard autoregressive decoding. In speculative decoding, let $T_{D}$ be the draft model’s per-token generation time, and let $T_{V}(\gamma)$ be the time for the target to verify $\gamma$ proposed tokens. Given $\alpha\in[0,1]$ as the expected acceptance rate and $\gamma\in\mathbb{N}$ the speculation length, following standard derivations (leviathan2023fast), the expected number of tokens produced per verification step is
@@ 3.2 Impact of Static Sub-Vocabulary Constraints on Acceptance
 [^1] For details, follow section 3 in (leviathan2023fast)
@@
$\beta(c;q_{c})$ $=\;\sum_{x\in V}\min(p_{c}(x),\,q_{c}(x))$ (3)
$=\;\mathbb{E}_{x\sim q_{c}}\!\left[\min\!\Big(1,\tfrac{p_{c}(x)}{q_{c}(x)}\Big)\right].$ (4)
@@
$A^{\!*}(c;S)$ $:=\;\max_{q_{c}:\ \mathrm{supp}(q_{c})\subseteq S}\ \sum_{x\in V}\min\{p_{c}(x),q_{c}(x)\}$
$=\;\sum_{x\in S}p_{c}(x)\,.$
@@
$\mathbb{E}_{c}\!\big[A^{\!*}(c;S_{c}^{\star})\big]=\;\mathbb{E}_{c}\!\big[H_{k}(p_{c})\big]\;\geq\;\mathbb{E}_{c}\!\big[A^{\!*}(c;S_{\mathrm{stat}})\big]$
$=\;\mathbb{E}_{c}\!\Big[\sum_{x\in S_{\mathrm{stat}}}p_{c}(x)\Big],$
@@ 4.1 Preliminary
Full Vocabulary Drafter. Speculative decoding methods (more specifically EAGLE (li2024eagle; li2024eagle2) series of models) train a lightweight draft model $\mathcal{D}$ (one transformer block, same latent width $d$) to approximate $\mathcal{T}$ at a lower cost. $\mathcal{D}$ reuses and freezes the embedding $\mathbf{E}$ and LM head $W_{LM}$ from $\mathcal{T}$; only its single transformer block is trained:
@@
Static Sub-vocabulary Drafter. FR-Spec (zhao-etal-2025-fr) uses a frequency-ranked drafting with vocabulary $\mathcal{V}_{\text{high}}\!\subset\!\mathcal{V}$ the subset of high-frequency tokens (from corpus statistics). Concretely, select the corresponding LM-head columns.
@@ 4.2 Meta classifiers: Context-aware token Selection
Instead of a fixed top-$k$ policy, we use a position-aware budget $k_{\mathrm{c}}(t)\in\{k_{\min},\ldots,k_{\max}\}$ that allocates more clusters to early tokens and fewer to later ones. Concretely, with $t=0$ denoting the first drafted token, we assign a larger initial budget and then decay as
$k_{\mathrm{c}}(t)=\begin{cases}k_{0},&t=0,\\[4.0pt] k_{1},&t=1,\\[4.0pt] \left\lfloor\dfrac{k_{1}+t}{2t}\right\rfloor,&t\geq 2,\end{cases}$
where $k_{0}>k_{1}$ and both are hyperparameters controlling the initial cluster budgets. Given router scores $s$, we select $\mathcal{K}(c,t)=\operatorname{TopK}_{k_{\mathrm{c}}(t)}(s)$ and construct the shortlist $\mathcal{V}_{S}(c,t)=\bigcup_{m\in\mathcal{K}(c,t)}C_{m}$.
@@
$\mathcal{L}_{\mathrm{r}}\;=\;\mathbb{E}\!\left[\sum_{m=1}^{M}\Big(-\,y_{m}\log\hat{y}_{m}-(1-y_{m})\log\big(1-\hat{y}_{m}\big)\Big)\right]$
Further training details are provided in A.2.
Figure 2: Time breakdown of the drafting process of DynaSpec. $T_{\text{emb}}$ is the embedding time, $T_{\text{core}}$ is the encoder layer time, $T_{\text{meta}}$ is the time related to the meta classifier and $T_{\text{index+gemm}}$ is the indexed matmul time.
4.3 Dynamic Drafting
The drafting process is depicted in Figure 1. At each step $t$, the router (meta-classifier) selects the shortlist indices, while the drafter generates the hidden state for that step. A single gathered-head GEMM is then performed over the shortlisted tokens using this hidden state. The router runs on a separate CUDA stream and typically finishes before the drafter block as shown in Figure 2. The total draft time for a single pass can be expressed as,
$T_{D}(c,t)\ \approx\ T_{\text{emb}}+\max\!\big\{T_{\text{core}},\,T_{\text{meta}}\big\}\ +T_{\text{index+gemm}}\!\big(B(c,t)\big)$
where $T_{\text{meta}}\!\ll\!T_{\text{core}}$ and $T_{\text{emb}}$ is generally small. Thus the only vocabulary-dependent term is the $d\times B(c,t)$ head compute (with $B(c,t)\!\ll\!|\mathcal{V}|$), reducing the denominator in Eq. 2 while preserving acceptance via context-aware support. Since $B(c,t)\!\propto\!|\mathcal{V}_{S}(c,t)|$ and the position-aware budget $k_{\mathrm{c}}(t)$ decay with $t$, the shortlist and thus the head compute shrink as generation proceeds, while larger early budgets preserve acceptance. In our implementation, index selection and the head GEMM are fused into a single custom CUDA kernel to avoid intermediate buffers and redundant global-memory traffic: the kernel performs warp-level column gathers with asynchronous, double-buffered shared-memory tiling, computes on Tensor Cores, and uses a persistent L2 access-policy window to keep activations cached. Verification is unchanged from standard speculative decoding.
5 Experimental Results
5.1 Experimental Settings
We adhere to the experimental setup and evaluation of the EAGLE (li2024eagle; li2024eagle2) family and FR-Spec (zhao-etal-2025-fr). We implement the DynaSpec in the FR-Spec  [^2] https://github.com/thunlp/FR-Spec framework. All components are implemented in native C++ and CUDA with FR-Spec as the baseline, since Python call overhead would otherwise negate the latency gains from draft-time acceleration.
Datasets and Models. We benchmark DynaSpec over a diverse mix of seven tasks. Spec-Bench (xia2024unlocking) supplies six of them—Machine Translation (WMT14 DE to EN) (bojar2014findings), Multi-turn Conversation (MT-Bench) (zheng2023eric), RAG and QA from Natural Questions (kwiatkowski2019natural), Math from GSM8K (cobbe2021training), and Summarization (CNN/DailyMail) (nallapati2016abstractive)—each evaluated on 80 prompts. The seventh task is Code, for which we employ HumanEval (chen2021evaluating) with 164 problems. We bound generations to $1024$ tokens for Spec-Bench subtasks and HumanEval. We used Llama series of models, specifically Llama-33-8B-Instruct, which uses a vocabulary size of $128K$ for our experiments and Qwen-2-7B.
Evaluation metrics and Hardware Settings. We evaluate using metrics: #Mean Acceptance Length, the average number of tokens committed per draft–verification cycle. #Mean Acceptance Length is hardware-agnostic but does not capture the overhead of drafting. Our primary baseline is FR-Spec (zhao-etal-2025-fr), a static subset–selection approach built on EAGLE-2 (li2024eagle2); in contrast, DynaSpec performs dynamic subset selection atop EAGLE-2 (li2024eagle2). All experiments were run on a single NVIDIA A6000 GPU.
Model / Config MT Conv. RAG Math QA Summ. Code Average Llama-3-8B Full Vocab $3.65$ $4.13$ $4.04$ $4.32$ $3.46$ $3.69$ $3.92$ $3.89$ vocab size=128k FR-Spec $3.39$ $3.88$ $3.86$ $4.16$ $3.33$ $3.52$ $3.31$ $3.64$ vocab size=32k DynaSpec $3.51$ $4.07$ $3.93$ $4.23$ $3.38$ $3.55$ $3.85$ $3.79$ vocab size 27,477 26,989 27,358 27,210 27,019 26,933 26,168 27,022 Qwen-2-7B Full Vocab $2.92$ $3.79$ $3.43$ $4.34$ $3.07$ $3.43$ $4.17$ $3.59$ vocab size=152k FR-Spec $2.84$ $3.52$ $3.30$ $4.17$ $2.95$ $3.32$ $3.52$ $3.37$ vocab size=32k DynaSpec $2.86$ $3.72$ $3.32$ $4.18$ $2.97$ $3.24$ $3.96$ $3.46$ vocab size=27k
Table 1: Mean accepted tokens for Llama-3-8B and Qwen-2-7B across seven datasets. vocab size indicates the shortlisted number of tokens. Full Vocab corresponds to the EAGLE-2 baseline; FR-Spec uses 32k. For DynaSpec, shortlist size varies by dataset; per-dataset means are shown individually.
Method MT. Conv. RAG Math QA Summ. Code Average
[…]
Table 2: Average tokens/sec for Llama-33-8B-Instruct on seven datasets. Numbers in parentheses under Average indicate the ratio vs. the Target baseline.